For the second question please.1. Given G(a) = +2 sketch the Nyquist plot and determine the range of K for closed-loop stability. 2. Repeat for G(s) = (+1)For the second question please.
Answer #1

Polca / Nyquist plot: It is a plot of magnitude in polae coddinater while w and es phase varied of GCS) H(S) from o to co.From the above ear, it is seen that G (jw) may be represented as a phasor of magnitude M and phase angle & As the ill frequenNow Given problem is Q- TK 1 968) ?2) Gls) = G(S) H() = k S(5+1)2 S(5+12 we to need magnitude and phase. to M find La at w=0 - 90 - tain (W) - tann (w . اہم کرNow at w=wp MLP At w= wpc Law) H(o) = -185 - - 180 = -98 - tant(6) - tanille) ao =. 2 tan I w tan (w) = 45 w = wpc = 1Altanate method: (Crest check ) R-H citeria Chae an is 33 +25² + stk L -2 For be a any system sign to be changes stable in th